# On the Asymptotic Mean Square Error Optimality of Diffusion Models

### v2

@@ 1 Introduction
Diffusion [1, 2] and score-based models [3, 4] are the backbone of foundation models such as Stable Diffusion [5]. Their success over different generative models, e.g., generative adversarial networks (GANs) [6] or variational autoencoders (VAEs) [7], is built upon their improved generative ability [8] and their theoretical understanding, particularly in terms of the convergence to the prior distribution of deterministic-time [9, 10] and continuous-time DMs [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21] with polynomial-time convergence guarantees. This advancement has led to the design of novel algorithms that utilize DMs, e.g., for image denoising [22, 23, 24], inverse problems [25, 26], super-resolution [27], and image restoration [28]. Moreover, DMs as generative priors have been successfully applied in various domains [29, 30, 31, 32].
@@
When considering denoising, the CME is the Bayesian-optimal solution in terms of the MSE [33], which is closely related to the peak signal-to-noise ratio (PSNR). It is of great importance in estimation and information theory [34, 35], statistical signal processing [36], and a pivotal benchmark for the analysis of denoising problems [37]. Through Tweedie’s formula [38], the CME can be directly related to the score, which was, e.g., exploited in Noise2Score [39]. Moreover, the minimum mean square error (MMSE) denoiser is frequently utilized in plug-and-play approaches [40, 41] for which DMs are particularly viable [28]. In image processing or related fields, the MMSE is an imperative metric to assess the perception-distortion trade-off [37, 42, 43]. In addition, in applications where the perceptual quality is not directly measurable, the CME is a desirable performance bound, e.g., in speech enhancement [44], wireless communications [45, 46], biomedical applications [47], or for natural signals in the wavelet domain [48].
@@
Despite the prevailing usage of DMs in estimation tasks [26, 27, 49, 50, 51], the solid theoretical guarantees for the convergence to the prior distribution [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21], and the importance of the CME in various fields, there is a notable gap in the theoretical understanding of the connection of DM-based denoising and the CME. In view of the pivotal role of DMs for foundational models and their excellent denoising capabilities, a natural question to ask is: How can a pre-trained DM be utilized in a fast and efficient manner for MSE-optimal estimation with provable convergence guarantees?
As shown in this paper, convergence to the prior distribution alone is insufficient to conclude about the convergence of a diffusion-based denoiser to the MSE-optimal CME. Thus, we contribute novel insights by proposing to leverage a pre-trained DM to deploy a fast and efficient denoising strategy with provable polynomial-time error guarantees. In short, we show the following main result.
@@
Let ${\bm{y}}={\bm{x}}+{\bm{n}}\in\mathbb{R}^{N}$ be a given observation that is corrupted by additive white Gaussian noise (AWGN) ${\bm{n}}$ and the stepwise error of the DM estimate be bounded by $\Delta$. Then, the distance of the proposed denoiser $f_{{\bm{\theta}}}$ that is parameterized through a pre-trained DM with $T$ timesteps to the MSE-optimal CME $\mathbb{E}[{\bm{x}}|{\bm{y}}]$ is bounded as
@@
  1. 1.
    We motivate a novel denoising strategy that utilizes a pre-trained DM but only forwards the stepwise conditional mean in the inference phase without drawing stochastic samples in the reverse process by showing a connection to the ground-truth CME. This results in a fast and efficient inference procedure that lays the foundation for the theoretical analysis.
  2. 2.
    We derive a new closed-form expression of the DM’s Lipschitz constant solely depending on the DM’s hyperparameters, which holds under mild assumptions; in particular, holds regardless of the stepwise estimation error of the DM.
  3. 3.
    We connect the optimality analysis of the proposed DM-based denoiser to the convergence analysis of the prior distribution, revealing that even uniform convergence to the prior distribution is insufficient to directly conclude about the convergence to the MSE-optimal CME, highlighting the importance of our theoretical analysis.
  4. 4.
    In our main result, we rigorously prove that the proposed DM-based denoising strategy is close to the ground-truth CME with polynomial-time convergence guarantee based on the derived Lipschitz constant under mild assumptions, i.e., we only require the common assumption of a bounded stepwise error of the DM, without assumptions about the convergence to the prior distribution or otherwise restrictive assumptions about the data distribution.
  5. 5.
    We reveal a novel perspective that DMs are comprised of a powerful generative model and an asymptotically MSE-optimal denoiser at the same time by simply switching the stochastic re-sampling in the reverse process on and off, respectively. This is an important contribution to the research of utilizing DMs as foundation models.
  6. 6.
    We thoroughly validate the theoretical findings by experiments based on various benchmark datasets, demonstrating that the proposed DM-based denoising strategy is capable of achieving the MMSE performance. Ablation studies based on a mismatch in the observation’s SNR information further highlight the remarkable robustness of the approach.
@@ 2 Related Work
Deterministic Sampling with DMs. Designing a DM with a deterministic reverse process is of great interest in current research since it allows for accelerated sampling and simplifies the design of a model consistent with the data manifold. However, building a DM with the mentioned properties necessitates a sophisticated training procedure since sampling from a fixed point in the reverse process does not sample the whole distribution [52]. The most prominent work investigating deterministic sampling is [53], where a non-Markovian forward process is designed to lead to a Markovian reverse process. This allows for a deterministic generative process where no re-sampling is necessary in each step. In [54], a consistency model is designed with a one-step network that maps a noise sample to its corresponding datapoint. The difference to our work is that the goal in these works is generation rather than denoising, and no pre-trained model can be used.
@@
Diffusion-based Denoising Approaches Denoising has a large history in machine learning [37], and utilizing powerful generative priors for denoising tasks is a well-known and promising strategy [55]. Building on DMs, stochastic denoising, i.e., posterior sampling, was proposed [27, 26, 49, 50, 51]. Alternatively, conditional DMs were designed where the DM network is trained conditioned on a noisy version of the true sample [27, 49]. The authors of [22, 23] propose a new denoising strategy by starting the reverse sampling at an intermediate timestep that corresponds to the noise level of the observation rather than from pure noise to reduce the sampling time. This procedure is possible if the DM is consistent with the noise model. Earlier works studying sub-sampling of DMs in the context of generation rather than denoising to reduce sampling time are found in [56, 57]. Our work utilizes a similar strategy, i.e., the reverse sampling is started from the noise level of the observation using a pre-trained DM since we assume that the observations are corrupted with AWGN. However, we afterward use a deterministic reverse process and study the convergence to the CME rather than employing stochastic denoising. Another related work that was recently devised, sharing a similar idea of forwarding the conditional mean in each step, is inversion by direct iteration [58]. A fundamental difference is that they do not train a generative model but rather a nested regression process, making a variational optimization approach superfluous, entailing an entirely different convergence behavior.
@@
In contempt of these advancements, the connection between the DM’s inference process and its connection to the MSE-optimal CME remains unexplored, and the role of the stochastic re-sampling in the reverse process is not fully understood. Beyond that, an effective inference process to achieve the MSE-optimal solution utilizing a pre-trained DM is absent, and no theoretical results in this context are known, being a notable gap in the study of DMs that is addressed in this work.
@@
Convergence to the Prior Distribution. Many recent works have been devoted to the analysis of the DM’s convergence to the prior distribution under various conditions on the data distribution and technical assumptions of the DM, resulting in an extensive collection of different convergence results [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. The works [9, 10] are most related to our setup since they also consider discrete-time DMs. In [59], it is shown that the DM objective can provably recover the ground-truth parameters of a Gaussian mixture model (GMM) distribution. This result is of particular interest in our context since the MSE-optimal CME can be computed in closed-form for a GMM distribution [45, 60]. Similar to our main result, the DM’s hyperparameters and the stepwise error are assumed to entail an asymptotic behavior over the number of diffusion steps in [10]. Theoretical justifications of this natural assumption of a bounded stepwise DM error can be further found in [12, 14, 20]. However, although these strong results about the convergence to the prior distribution exist, no result discusses the theoretical capability of DMs to approximate the MSE-optimal CME. As shown in Section 4, a direct conclusion about the convergence to the CME given a convergence to the prior distribution is not straightforward, necessitating a careful investigation of the error bound.
@@ 3 Preliminaries
3.1 Problem Formulation
We consider a denoising task where the true sample is corrupted by AWGN, yielding an observation
@@ 3.1 Problem Formulation
$g_{\text{MMSE}}({\bm{y}})=\operatorname*{arg\,min}_{g:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}}\mathbb{E}[\|{\bm{x}}-g({\bm{y}})\|_{2}^{2}]$ (3)
where $g_{\text{MMSE}}$ is known as the CME, computed as
$g_{\text{MMSE}}({\bm{y}})$ $=\mathbb{E}[{\bm{x}}|{\bm{y}}]=\int_{\mathbb{R}^{N}}{\bm{x}}p({\bm{x}}|{\bm{y}}){\operatorname{d}}{\bm{x}}=\int_{\mathbb{R}^{N}}{\bm{x}}\frac{q({\bm{y}}|{\bm{x}})p({\bm{x}})}{p({\bm{y}})}{\operatorname{d}}{\bm{x}}.$ (4)
However, for an unknown prior $p({\bm{x}})$, the CME is intractable to compute due to the high-dimensional integral, and reasonable approximations must be found. A naive approach is to parameterize a one-step regression neural network (NN) $g_{{{\bm{\theta}}}}$ trained on the MSE loss (3) to provide an estimate $g_{{{\bm{\theta}}}}({\bm{y}})\approx g_{\text{MMSE}}({\bm{y}})$ at the output of the NN [61, 62]. However, in contrast to a regression network, the proposed DM-based denoiser has several significant advantages:
  • One-step regression networks are “task-specific” and have to be trained from scratch on paired examples (commonly noisy and clean samples) [37]. In contrast, a pre-trained DM can be considered a “general purpose” denoiser leveraging powerful foundational DMs such as Stable Diffusion [5].
  • One-step networks are generally prone to generalization issues when being trained on a large range of noise levels, necessitating either re-training or a meticulous design of the architecture and loss function [63, 64]. In contrast, DMs are expected to generalize better as they are inherently trained on the full SNR range.
  • One-step networks are not very well understood theoretically, as the convergence to the CME is solely attributed to the universal approximation property [65]. With this paper, we aim to address this limitation by rigorously proving the (non-)asymptotic convergence of the proposed DM-based denoising strategy to the CME under mild assumptions.
3.2 Diffusion Models
We briefly review the formulations of DMs from [2]. Given a data distribution ${\bm{x}}_{0}\sim p({\bm{x}}_{0})$, the forward process which produces latents ${\bm{x}}_{1}$ through ${\bm{x}}_{T}$ by adding Gaussian noise at time $t$ with the hyperparameters $\alpha_{t}\in(0,1)$ for all $t=1,\dots,T$ is a Markov chain that is defined via the transition
$q({\bm{x}}_{t}|{\bm{x}}_{t-1})=\mathcal{N}\left({\bm{x}}_{t};\sqrt{\alpha_{t}}{\bm{x}}_{t-1},(1-\alpha_{t})\operatorname{\mathbf{I}}\right).$ (5)
Applying the reparameterization trick (iteratively) lets us write
${\bm{x}}_{t}$ $=\sqrt{\alpha_{t}}{\bm{x}}_{t-1}+\sqrt{1-\alpha_{t}}{\bm{\epsilon}}_{t-1}=\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}_{0}$ (6)
@@ 3.2 Diffusion Models
$p_{{\bm{\theta}}}({\bm{x}}_{t-1}|{\bm{x}}_{t})=\mathcal{N}({\bm{x}}_{t-1};{\bm{\mu}}_{{{\bm{\theta}}}}({\bm{x}}_{t},t),{\bm{\Sigma}}_{{\bm{\theta}}}({\bm{x}}_{t},t)=\sigma_{t}^{2}\operatorname{\mathbf{I}}).$ (7)
As in [2], we set the variances of the reverse process transitions to untrained time-dependent constants. We note that the variances can also be chosen as learnable parameters, cf. [56, 66], which is not considered in this work. The choice in (7) is motivated by the fact that the forward and reverse process of a Markov chain have the same functional form when $\alpha_{t}$ is close to one for all $t=1,\dots,T$ [67, 1]. However, the transitions in (7) are generally intractable to compute and are thus learned via the forward posteriors, which are tractable when conditioned on ${\bm{x}}_{0}$, i.e.,
@@
$\tilde{{\bm{\mu}}}({\bm{x}}_{t},{\bm{x}}_{0})$ $=\frac{\sqrt{\bar{\alpha}_{t-1}}(1-\alpha_{t})}{1-\bar{\alpha}_{t}}{\bm{x}}_{0}+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}{\bm{x}}_{t},~{}~{}\sigma_{t}^{2}=\frac{(1-\alpha_{t})(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}.$ (9)
Details about the training of the DM through maximizing the evidence lower bound (ELBO) are given in Appendix A.
In the remainder of this work, we consider $\mathbb{E}[{\bm{x}}_{0}]={\bm{0}}$ and $\mathbb{E}[\|{\bm{x}}_{0}\|_{2}^{2}]=N$, which is ensured by a pre-processing step, such that the observation’s SNR can be defined as $\text{SNR}({\bm{y}})=1/\eta^{2}$. Then, the DM timesteps can be equivalently interpreted as different SNR steps, cf. [66, 68], by defining the DM’s SNR of step $t$, which is monotonically decreasing for increasing $t$, cf. (6), as
@@
4 MSE-Optimal Denoising with Diffusion Models
4.1 Proposed Denoising Procedure and Connection to the CME
Proposed Denoiser. We consider a DM that is consistent with the noise model, i.e., the latent variables in the forward process and the observation (2) are corrupted by AWGN. In that case, it is possible to initialize the reverse process at timestep ${\hat{t}}<T$ that corresponds to the SNR of the observation rather than from pure noise [22, 23]. This results in a considerable reduction of reverse steps for denoising, effectively reducing the computational complexity and allowing the utilization of a pre-trained DM. Thus, after normalizing the observation from (2), i.e., $\tilde{{\bm{y}}}=(1+\eta^{2})^{-\frac{1}{2}}{\bm{y}}$ such that $\mathbb{E}[\|\tilde{{\bm{y}}}\|^{2}_{2}]=N$, which is necessary since the DM is variance-preserving, we find the timestep of the DM that best matches the SNR of the observation, which is assumed to be known (we refer to Section 5 and Appendix K for an ablation study of a mismatch in the SNR information), via
${\hat{t}}=\operatorname*{arg\,min}_{t}|\text{SNR}({\bm{y}})-\text{SNR}_{\text{DM}}(t)|=\operatorname*{arg\,min}_{t}\left|\frac{1}{\eta^{2}}-\frac{\bar{\alpha}_{t}}{1-\bar{\alpha}_{t}}\right|.$ (11)
We note that ${\hat{t}}<T$ solely depends on the hyperparameter choice of the DM and can be either computed analytically or by a simple search over the DM’s timesteps. Consequently, we initialize the reverse process with ${\bm{x}}_{{\hat{t}}}=\tilde{{\bm{y}}}$ and denote the ground-truth CME as $\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{\hat{t}}]$. The CME is a deterministic point estimate, motivating to avoid re-sampling in the reverse process during inference if convergence to the CME is desired, as opposed to stochastic denoising [49, 23, 27, 26]. By denoting $f_{{{\bm{\theta}}},t}({\bm{x}}_{t}):={\bm{\mu}}_{{\bm{\theta}}}({\bm{x}}_{t},t)$, cf. (7), the proposed DM-based denoiser is given as

### v4

@@ 1 INTRODUCTION
DMs (Sohl-Dickstein et al.,, 2015; Ho et al.,, 2020) and score-based models (Song and Ermon,, 2019; Song et al., 2021b, ) are the backbone of foundation models such as Stable Diffusion (Rombach et al.,, 2022). Their success over different generative models, e.g., generative adversarial networks (GANs) (Goodfellow et al.,, 2014) or variational autoencoders (VAEs) (Kingma and Welling,, 2014), is built upon their improved generative ability (Dhariwal and Nichol,, 2021) and their theoretical understanding, particularly in terms of the convergence to the prior distribution of deterministic-time (Li et al.,, 2024; Liang et al.,, 2025) and continuous-time DMs (Bortoli et al.,, 2021; Chen et al., 2023c, ; Lee et al.,, 2023; Chen et al., 2023a, ; Benton et al.,, 2024; Pedrotti et al.,, 2024; Conforti et al.,, 2025; Chen et al., 2023b, ; Bortoli,, 2022; Block et al.,, 2022; Chen et al., 2023d, ) with polynomial-time convergence guarantees. This advancement has led to the design of novel algorithms that utilize DMs, e.g., for inverse problems (Chung et al.,, 2022; Meng and Kabashima,, 2024) and image denoising (Carlini et al.,, 2023; Xie et al.,, 2023; Fabian et al.,, 2024), super-resolution (Saharia et al.,, 2023), and restoration (Zhu et al.,, 2023). Moreover, DMs as generative priors have been successfully applied in various domains (Yang et al.,, 2023; Kong et al.,, 2021; Arvinte and Tamir,, 2023; Huy and Minh Quan,, 2023).
@@
When considering denoising, the CME is the Bayesian-optimal solution in terms of the MSE (Lehmann and Casella,, 2006), which is closely related to the peak signal-to-noise ratio (PSNR). It is of great importance in estimation and information theory (Kay,, 1993; Asoodeh et al.,, 2016), statistical signal processing (Scharf and Demeure,, 1991), and a pivotal benchmark for the analysis of denoising problems (Elad et al.,, 2023). Through Tweedie’s formula (Efron,, 2011), the CME can be directly related to the score, which was, e.g., exploited in Noise2Score (Kim and Ye,, 2021). Moreover, the minimum mean square error (MMSE) denoiser is frequently utilized in plug-and-play approaches (Venkatakrishnan et al.,, 2013; Romano et al.,, 2017) for which DMs are particularly viable (Zhu et al.,, 2023). In image processing or related fields, the MMSE is an imperative metric to assess the perception-distortion trade-off (Elad et al.,, 2023; Blau and Michaeli,, 2018; Om and Biswas,, 2014), which states that optimizing for perceptual quality often conflicts with minimizing distortion measures, e.g., the MSE. In addition, in applications where the perceptual quality is not directly measurable, the CME is a desirable performance bound, e.g., in speech enhancement (Loizou,, 2013), wireless communications (Koller et al.,, 2022; Fesl et al.,, 2023), biomedical applications (Gargiulo and McEwan,, 2011), or wavelet domain processing (Kazerouni et al.,, 2013).
@@
Despite the prevailing usage of DMs in estimation tasks (Meng and Kabashima,, 2024; Saharia et al.,, 2023; Whang et al.,, 2022; Huang et al.,, 2024; Tai et al.,, 2023), the solid theoretical guarantees for the convergence to the prior distribution (Li et al.,, 2024; Liang et al.,, 2025; Bortoli et al.,, 2021; Chen et al., 2023c, ; Lee et al.,, 2023; Chen et al., 2023a, ; Benton et al.,, 2024; Pedrotti et al.,, 2024; Conforti et al.,, 2025; Chen et al., 2023b, ; Bortoli,, 2022; Block et al.,, 2022; Chen et al., 2023d, ), and the importance of the CME in various fields, there is a notable gap in the theoretical understanding of the connection of DM-based denoising and the CME. In view of the pivotal role of DMs for foundational models and their excellent denoising capabilities, a natural question to ask is: How can a pre-trained DM be utilized in a fast and efficient manner for MSE-optimal estimation with provable convergence guarantees? In short, we show the following main result.
@@
Let ${\bm{y}}={\bm{x}}+{\bm{n}}\in\mathbb{R}^{N}$ be a noisy observation with additive white Gaussian noise (AWGN) ${\bm{n}}$ and the stepwise denoising error of the DM’s reverse process be bounded by $\Delta$. Then, the distance of the proposed denoiser $f_{{\bm{\theta}}}({\bm{y}})$ that utilizes a pre-trained DM with $T$ timesteps to the MSE-optimal CME $\mathbb{E}[{\bm{x}}|{\bm{y}}]$ is bounded as
@@
The error bound in (1) depends solely on the number of discretization steps $T$ and the bounded stepwise denoising error $\Delta$ of the DM. Moreover, the error bound approaches zero in polynomial time as the number of diffusion steps $T$ increases, making the proposed DM-based denoiser asymptotically optimal.
@@
1) We motivate a novel denoising strategy that utilizes a pre-trained DM but only forwards the stepwise conditional mean in the inference phase without drawing stochastic samples in the reverse process by showing a connection to the ground-truth CME. This results in a fast and efficient inference procedure that lays the foundation for the theoretical analysis.
2) We derive a new closed-form expression for the DM’s stepwise Lipschitz constant, which depends solely on the DM’s hyperparameters and holds under mild assumptions; in particular, it remains valid regardless of the stepwise estimation error of the DM. This novel Lipschitz constant is pivotal in deriving the (non-)asymptotic convergence results. Additionally, it serves as the foundation to demonstrate that the proposed DM-based denoiser inherently optimizes the bias-variance trade-off, a fundamental principle in estimation theory.
3) We connect the optimality analysis of the proposed DM-based denoiser to the convergence analysis of the prior distribution. In our main result, we rigorously prove that the proposed DM-based denoising strategy is close to the ground-truth CME with polynomial-time convergence guarantee based on the derived Lipschitz constant under mild assumptions, i.e., we only require the common assumption of a bounded stepwise error of the DM, without assumptions about the convergence to the prior distribution or otherwise restrictive assumptions about the data distribution.
4) We reveal a novel perspective that DMs are comprised of a powerful generative model and an asymptotically MSE-optimal denoiser at the same time by switching the stochastic re-sampling in the reverse process on and off. In addition, we thoroughly validate the theoretical findings by experiments based on various benchmark datasets and perform ablation studies, demonstrating that the proposed DM-based denoising strategy is robust and asymptotically optimal.
@@ 2 RELATED WORK
Deterministic Sampling with DMs. Designing a DM with a deterministic reverse process is of great interest in current research since it allows for accelerated sampling and simplifies the design of a model consistent with the data manifold. The most prominent work investigating deterministic sampling is from Song et al., 2021a , where a non-Markovian forward process is designed to lead to a Markovian reverse process. This allows for a deterministic generative process where no re-sampling is necessary in each step. Song et al., (2023) designed a consistency model with a one-step network that maps a noise sample to its corresponding datapoint. In contrast, our work presents a deterministic adaptation of the DM architecture presented by Ho et al., (2020), focusing on denoising rather than generation and allowing the use of pre-trained DMs.
Song et al., 2021b were the first to interpret deterministic DMs as solutions to the corresponding probability flow ordinary differential equation (ODE). Building on this perspective, subsequent research has introduced new findings, e.g., (Lu et al.,, 2022; Chen et al., 2023b, ; Kim et al.,, 2024; Zhou et al.,, 2024). The ODE formulation of DMs can also be viewed as a neural ODE (Chen et al.,, 2018), highlighting its similarities with continuous normalizing flows (Papamakarios et al.,, 2021). Lipman et al., (2023) proposed an architecture that integrates both methodologies for image generation. While our work focuses on the DM architecture proposed by Ho et al., (2020), our theoretical results may be extendable to these alternative approaches.
Diffusion-based Denoising Approaches. Denoising has a large history in machine learning (Elad et al.,, 2023), and utilizing powerful generative priors for denoising tasks is a well-known and promising strategy (Kadkhodaie and Simoncelli,, 2021). Building on DMs, stochastic denoising, i.e., posterior sampling, was proposed (Saharia et al.,, 2023; Meng and Kabashima,, 2024; Whang et al.,, 2022; Huang et al.,, 2024; Tai et al.,, 2023).
@@
Alternatively, conditional DMs were designed where the DM network is trained conditioned on a noisy version of the true sample (Saharia et al.,, 2023; Whang et al.,, 2022). Carlini et al., (2023) and Xie et al., (2023) proposed a new denoising strategy by starting the reverse sampling at an intermediate timestep that corresponds to the noise level of the observation rather than from pure noise to reduce the sampling time. This procedure is possible if the DM is consistent with the noise model. Nichol and Dhariwal, (2021) and Xiao et al., (2023) also studied sub-sampling of DMs in the context of generation rather than denoising to reduce sampling time. Our work utilizes a similar strategy, i.e., the reverse sampling is started from the noise level of the observation using a pre-trained DM since we assume that the observations are corrupted with AWGN. However, we afterward use a deterministic reverse process and study the convergence to the CME rather than employing stochastic denoising.
Another related work that was recently devised, sharing a similar idea of forwarding the conditional mean in each step, is from Delbracio and Milanfar, (2023). A fundamental difference is that they do not train a generative model but rather a nested regression process, entailing an entirely different convergence behavior.
@@
In contempt of these advancements, the connection between the DM’s inference process and the MSE-optimal CME remains unexplored, and the role of the stochastic re-sampling in the reverse process is not fully understood. Beyond that, an effective inference process to achieve the MSE-optimal solution utilizing a pre-trained DM is absent, being a notable gap in the study of DMs that is addressed in this work.
@@
Convergence to the Prior Distribution. Many recent works have been devoted to the analysis of the DM’s convergence to the prior distribution under various conditions on the data distribution and technical assumptions of the DM, resulting in an extensive collection of different convergence results (Li et al.,, 2024; Liang et al.,, 2025; Bortoli et al.,, 2021; Chen et al., 2023c, ; Lee et al.,, 2023; Chen et al., 2023a, ; Benton et al.,, 2024; Pedrotti et al.,, 2024; Conforti et al.,, 2025; Chen et al., 2023b, ; Bortoli,, 2022; Block et al.,, 2022). The works from Li et al., (2024) and Liang et al., (2025) are most related to our setup since they also consider discrete-time DMs. Shah et al., (2023) showed that the DM can provably recover the ground-truth parameters of a Gaussian mixture model (GMM) distribution. This result is of particular interest in our context since the MSE-optimal CME can be computed in closed-form for a GMM distribution (Koller et al.,, 2022; Yang et al.,, 2015). Similar to our main result, the DM’s hyperparameters and the stepwise error are assumed to entail an asymptotic behavior over the number of diffusion steps by Liang et al., (2025). Several works discussed theoretical justifications of this natural assumption of a bounded stepwise DM error (Chen et al., 2023c, ; Chen et al., 2023a, ; Block et al.,, 2022). However, although these strong results about the convergence to the prior distribution exist, no result discusses the theoretical capability of DMs to approximate the MSE-optimal CME.
@@ 3 PRELIMINARIES
Problem Formulation. We consider a denoising task where the true sample is corrupted by AWGN, yielding an observation
@@
$g^{*}({\bm{y}})=\operatorname*{arg\,min}_{g:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}}\mathbb{E}[\|{\bm{x}}-g({\bm{y}})\|_{2}^{2}]$ (3)
where $g^{*}$ is known as the CME, computed as
$g^{*}({\bm{y}})$ $=\mathbb{E}[{\bm{x}}|{\bm{y}}]=\int{\bm{x}}p({\bm{x}}|{\bm{y}}){\operatorname{d}}{\bm{x}}=\int{\bm{x}}\frac{p({\bm{y}}|{\bm{x}})p({\bm{x}})}{p({\bm{y}})}{\operatorname{d}}{\bm{x}}.$ (4)
However, for an unknown prior $p({\bm{x}})$, the CME is intractable to compute due to the high-dimensional integral, and reasonable approximations must be found. A naive approach is to parameterize a one-step regression neural network (NN) $g_{{{\bm{\theta}}}}$ trained on the MSE loss (3) to provide an estimate $g_{{{\bm{\theta}}}}({\bm{y}})\approx g^{*}({\bm{y}})$ at the output of the NN (Dong et al.,, 2016; Ongie et al.,, 2020). However, in contrast to a regression network, the proposed DM-based denoiser has several significant advantages.
First, one-step regression networks are “task-specific” and have to be trained from scratch on paired examples (commonly noisy and clean samples) (Elad et al.,, 2023). In contrast, a pre-trained DM can be considered a “general purpose” denoiser leveraging powerful foundational DMs such as Stable Diffusion (Rombach et al.,, 2022). Second, one-step networks are generally prone to generalization issues when being trained on a large range of noise levels, necessitating either re-training or a meticulous design of the architecture and loss function (Zhao et al.,, 2017; Mohan et al.,, 2020). In contrast, DMs are expected to generalize better as they are inherently trained on the full SNR range. Third, one-step networks are not very well understood theoretically, as the convergence to the CME is solely attributed to the universal approximation property (Diaz et al.,, 2021). With this paper, we aim to address this limitation by rigorously proving the (non-)asymptotic convergence of the proposed DM-based denoising strategy to the CME under mild assumptions.
Diffusion Models. We briefly review the DM formulations from Ho et al., (2020). Given a data distribution ${\bm{x}}_{0}\sim p({\bm{x}}_{0})$, the forward process which produces latents ${\bm{x}}_{1}$ through ${\bm{x}}_{T}$ by adding Gaussian noise at time $t$ with the hyperparameters $\alpha_{t},\beta_{t}\in(0,1)$ with $\beta_{t}=1-\alpha_{t}$ for all $t=1,\dots,T$ is a Markov chain that is defined via the transition
$q({\bm{x}}_{t}|{\bm{x}}_{t-1})=\mathcal{N}\left({\bm{x}}_{t};\sqrt{\alpha_{t}}{\bm{x}}_{t-1},\beta_{t}\operatorname{\mathbf{I}}\right).$ (5)
Iteratively applying the reparameterization trick yields
${\bm{x}}_{t}$ $=\sqrt{\alpha_{t}}{\bm{x}}_{t-1}+\sqrt{\beta_{t}}{\bm{\epsilon}}_{t-1}=\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}_{0}$ (6)
@@
$p_{{\bm{\theta}}}({\bm{x}}_{t-1}|{\bm{x}}_{t})=\mathcal{N}({\bm{x}}_{t-1};{\bm{\mu}}_{{{\bm{\theta}}}}({\bm{x}}_{t},t),\sigma_{t}^{2}\operatorname{\mathbf{I}}).$ (7)
As Ho et al., (2020), we set the variances of the reverse process to untrained time-dependent constants. The choice in (7) is motivated by the fact that the forward and reverse process of a Markov chain have the same functional form when $\alpha_{t}$ is close to one for all $t=1,\dots,T$ (Feller,, 1949; Sohl-Dickstein et al.,, 2015). However, the transitions in (7) are generally intractable and are thus learned via the forward posteriors, which are tractable when conditioned on ${\bm{x}}_{0}$, i.e.,
@@
$\tilde{{\bm{\mu}}}({\bm{x}}_{t},{\bm{x}}_{0})$ $=\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}{\bm{x}}_{0}+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}{\bm{x}}_{t},$ (9)
$\sigma_{t}^{2}$ $=\frac{(1-\alpha_{t})(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}.$ (10)
Training details of the DM are given in Appendix A.
In the remainder of this work, we consider $\mathbb{E}[{\bm{x}}_{0}]={\bm{0}}$ and $\mathbb{E}[\|{\bm{x}}_{0}\|_{2}^{2}]=N$, which is ensured by a pre-processing step, such that the observation’s SNR can be defined as $\text{SNR}({\bm{y}})=1/\eta^{2}$. Then, the DM timesteps can be equivalently interpreted as different SNR steps (Kingma et al.,, 2021; Luo,, 2022) by defining the DM’s SNR of step $t$, which is monotonically decreasing for increasing $t$, cf. (6), as
@@
4 MSE-OPTIMAL DENOISING
4.1 Discussion of the DM-based Denoiser
Proposed Denoiser. We consider a DM that is consistent with the noise model, i.e., the latent variables in the forward process and the observation (2) are corrupted by AWGN. In that case, it is possible to initialize the reverse process at timestep ${\hat{t}}<T$ that corresponds to the SNR of the observation rather than from pure noise (Carlini et al.,, 2023; Xie et al.,, 2023). This results in a considerable reduction of reverse steps for denoising, effectively reducing the computational complexity and allowing the utilization of a pre-trained DM. Thus, after normalizing the observation from (2), i.e., $\tilde{{\bm{y}}}=(1+\eta^{2})^{-\frac{1}{2}}{\bm{y}}$ such that $\mathbb{E}[\|\tilde{{\bm{y}}}\|^{2}_{2}]=N$, which is necessary since the DM is variance-preserving, we find the timestep of the DM that best matches the SNR of the observation, which is assumed to be known (we refer to Section 5 and Appendix J for an ablation study of a mismatch in the SNR information), via
${\hat{t}}$ $=\operatorname*{arg\,min}_{t}|\text{SNR}({\bm{y}})-\text{SNR}_{\text{DM}}(t)|$ (12)
$=\operatorname*{arg\,min}_{t}\big{|}\frac{1}{\eta^{2}}-\frac{\bar{\alpha}_{t}}{1-\bar{\alpha}_{t}}\big{|}.$
We note that ${\hat{t}}<T$ solely depends on the hyperparameter choice of the DM and can be either computed analytically or by a simple search over the DM’s timesteps. Consequently, we initialize the reverse process with ${\bm{x}}_{{\hat{t}}}=\tilde{{\bm{y}}}$ and denote the ground-truth CME as $\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{\hat{t}}]$. The CME is a deterministic point estimate, motivating to avoid re-sampling in the reverse process during inference if convergence to the CME is desired, as opposed to stochastic denoising (Whang et al.,, 2022; Xie et al.,, 2023; Saharia et al.,, 2023; Meng and Kabashima,, 2024). Thus, instead of stochastic sampling from the parameterized reverse process distribution as ${\bm{x}}_{t-1}\sim\mathcal{N}({\bm{\mu}}_{{{\bm{\theta}}}}({\bm{x}}_{t},t),\sigma_{t}^{2}\operatorname{\mathbf{I}})$, cf. (7), we propose to solely forward its conditional mean ${\bm{\mu}}_{{{\bm{\theta}}}}({\bm{x}}_{t},t)$ in each step. By denoting $f_{{{\bm{\theta}}},t}({\bm{x}}_{t}):={\bm{\mu}}_{{\bm{\theta}}}({\bm{x}}_{t},t)$, cf. (7), the proposed DM-based denoiser is given as